# Multi-objective Reinforcement Learning with Nonlinear Preferences: Provable Approximation for Maximizing Expected Scalarized Return

### v3

@@ Appendix C Details on the RAEE Algorithm
In this section, we give a detailed description of the RAEE algorithm, extending from [16] and demonstrating Theorem 2, the main theorem for the RAEE algorithm.
C.1 Introducing RAEE
@@
The extension works as follows. The algorithm starts off by doing balanced wandering [16]: when encountering a new state, the algorithm selects a random action. However, when revisiting a previously visited state, it chooses the least attempted action from that state, resolving ties by random action selection. At each state-action pair $(s,a)$ it tries, the algorithm stores the reward $\mathbf{R}(s,a)$ received and an estimate of the transition probabilities $Pr(s^{\prime}\mid s,a)$ derived from the empirical distribution of next states reached during balanced wandering.
@@
Next, we introduce the notion of a known state [16], which refers to a state that the algorithm has explored to the extent that the estimated transition probabilities for any action from that state closely approximate their actual values. Denote the number of times a state needs to be visited as $m_{known}$. We will specify the value later in our runtime characterization.
@@
States are thus categorized into three groups: known states, which the algorithm has extensively visited and obtained reliable transition statistics; states that have been visited before but remain unknown due to limited trials and therefore unreliable data; and states that have not been explored at all. By the Pigeonhole Principle, accurate statistics will eventually accumulate in some states over time, leading to their becoming known. Let $S$ be the set of currently known states, the algorithm can build the current known-state MOMDP $M_{S}$ that is naturally induced on $S$ by the full MOMDP $M$ with all “unknown” states merged into a single absorbing state. Although the algorithm cannot access $M_{S}$ directly, it will have an approximation $\hat{M}_{S}$ by the definition of the known states. By the simulation lemma [16], $\hat{M}_{S}$ will be an accurate model in the sense that the expected $T$-step welfare of any policy in $\hat{M}_{S}$ is close to its expected $T$-step return in $M_{S}$. (Here $T$ is the horizon time.) Hence, at any timestep, $\hat{M}_{S}$ functions as an incomplete representation of $M$, specifically focusing on the part of $M$ that the algorithm possesses a strong understanding of.
@@
By the explore or exploit lemma [16], the algorithm is guaranteed to either output a policy with approximately optimal return in $M$, or to improve the statistics at an unknown state. Again by the Pigeonhole Principle, a new state becomes known after the latter case occurs for some finite number of times, and thus the algorithm is always making progress. In the worst case, the algorithm builds a model of the entire MOMDP $M$. Having described the elements, we now outline the entire extended algorithm, where the notations are consistent with [16].
@@
  • (Discovery of New Known States) Any time a state $s$ has been visited $m_{known}$ [16] times during balanced wandering, it enters the known set $S$, and no longer participates in balanced wandering.
@@ C.2 Runtime Analysis
[16] Let $M$ and $M^{\prime}$ be two MOMDPs over the same state space $\mathcal{S}$ with the same deterministic reward function $\mathbf{R}(s,a)$. $M^{\prime}$ is an $\alpha-$approximation of $M$ if for any state $s$ and $s^{\prime}$ and any action $a$,
@@
We now extend the Simulation Lemma [16], which tells us how close the approximation of an MOMDP needs to be in order for the expected welfare, or ESR, of a policy to be close in an estimated model. The argument is similar to [16], but in our case the policy may not be stationary and we want to bound the deviation in expected welfare rather than just accumulated reward.
@@
Let $M^{\prime}$ be an $O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$-approximation of $M$. Then for any policy $\pi$, any state $s$, and horizon time $T$, we have
@@
[16] Let $M$ be an MOMDP. A state $s$ of $M$ is considered known if it has be visited a number of times equal to
@@
[16] Let $M$ be an MOMDP. Let $s$ be a state of $M$ that has been visited at least $m$ times, with each action having been executed at least $\lfloor m/|\mathcal{A}|\rfloor$ times. Let $\hat{Pr}(s^{\prime}\mid s,a)$ denote the empirical transition probability estimates obtained from the $m$ visits to $s$. Then if
@@
$\bar{X}\in[p-\alpha,p+\alpha]~{}\text{with probability $1-\delta$}.$
@@
We have specified the degree of approximation required for sufficient simulation accuracy. It remains to directly apply the Explore or Exploit Lemma from [16] to conclude the analysis.
Lemma 7 (Explore or Exploit Lemma [16]).
@@
It remains to handle the failure parameter $\beta$ in the statement of the main theorem, which can be done similarly to [16]. There are two sources of failure for the algorithm:

### v4

@@ Appendix D Details on the RAEE Algorithm
In this section, we give a detailed description of the RAEE algorithm, extending from Kearns and Singh (2002) and demonstrating Theorem 12, the main theorem for the RAEE algorithm.
D.1. Introducing RAEE
@@
The extension works as follows. The algorithm starts off by doing balanced wandering Kearns and Singh (2002): when encountering a new state, the algorithm selects a random action. However, when revisiting a previously visited state, it chooses the least attempted action from that state, resolving ties by random action selection. At each state-action pair $(s,a)$ it tries, the algorithm stores the reward $\mathbf{R}(s,a)$ received and an estimate of the transition probabilities $Pr(s^{\prime}\mid s,a)$ derived from the empirical distribution of next states reached during balanced wandering.
@@
Next, we introduce the notion of a known state Kearns and Singh (2002), which refers to a state that the algorithm has explored to the extent that the estimated transition probabilities for any action from that state closely approximate their actual values. Denote the number of times a state needs to be visited as $m_{known}$. We will specify the value later in our runtime characterization.
@@
States are thus categorized into three groups: known states, which the algorithm has extensively visited and obtained reliable transition statistics; states that have been visited before but remain unknown due to limited trials and therefore unreliable data; and states that have not been explored at all. By the Pigeonhole Principle, accurate statistics will eventually accumulate in some states over time, leading to their becoming known. Let $S$ be the set of currently known states, the algorithm can build the current known-state MOMDP $M_{S}$ that is naturally induced on $S$ by the full MOMDP $M$ with all “unknown” states merged into a single absorbing state. Although the algorithm cannot access $M_{S}$ directly, it will have an approximation $\hat{M}_{S}$ by the definition of the known states. By the simulation lemma Kearns and Singh (2002), $\hat{M}_{S}$ will be an accurate model in the sense that the expected $T$-step welfare of any policy in $\hat{M}_{S}$ is close to its expected $T$-step return in $M_{S}$. (Here $T$ is the horizon time.) Hence, at any timestep, $\hat{M}_{S}$ functions as an incomplete representation of $M$, specifically focusing on the part of $M$ that the algorithm possesses a strong understanding of.
@@
By the explore or exploit lemma Kearns and Singh (2002), the algorithm is guaranteed to either output a policy with approximately optimal return in $M$, or to improve the statistics at an unknown state. Again by the Pigeonhole Principle, a new state becomes known after the latter case occurs for some finite number of times, and thus the algorithm is always making progress. In the worst case, the algorithm builds a model of the entire MOMDP $M$. Having described the elements, we now outline the entire extended algorithm, where the notations are consistent with Kearns and Singh (2002).
@@
  • (Discovery of New Known States) Any time a state $s$ has been visited $m_{known}$ Kearns and Singh (2002) times during balanced wandering, it enters the known set $S$, and no longer participates in balanced wandering.
@@ D.2. Runtime Analysis
Kearns and Singh (2002) Let $M$ and $M^{\prime}$ be two MOMDPs over the same state space $\mathcal{S}$ with the same deterministic reward function $\mathbf{R}(s,a)$. $M^{\prime}$ is an $\alpha-$approximation of $M$ if for any state $s$ and $s^{\prime}$ and any action $a$,
@@
We now extend the Simulation Lemma Kearns and Singh (2002), which tells us how close the approximation of an MOMDP needs to be in order for the expected welfare, or ESR, of a policy to be close in an estimated model. The argument is similar to Kearns and Singh (2002), but in our case the policy may not be stationary and we want to bound the deviation in expected welfare rather than just accumulated reward.
@@
Let $M^{\prime}$ be an
$O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$-approximation of $M$. Then for any policy $\pi$, any state $s$, and horizon time $T$, we have
@@
Kearns and Singh (2002) Let $M$ be an MOMDP. A state $s$ of $M$ is considered known if it has be visited a number of times equal to
@@
Kearns and Singh (2002) Let $M$ be an MOMDP. Let $s$ be a state of $M$ that has been visited at least $m$ times, with each action having been executed at least $\lfloor m/|\mathcal{A}|\rfloor$ times. Let $\hat{Pr}(s^{\prime}\mid s,a)$ denote the empirical transition probability estimates obtained from the $m$ visits to $s$. Then if
@@
$\bar{X}\in[p-\alpha,p+\alpha]\leavevmode\nobreak\ \text{with probability $1-\delta$}.$
@@
We have specified the degree of approximation required for sufficient simulation accuracy. It remains to directly apply the Explore or Exploit Lemma from Kearns and Singh (2002) to conclude the analysis.
Lemma 0 (Explore or Exploit Lemma Kearns and Singh (2002)).
@@
It remains to handle the failure parameter $\beta$ in the statement of the main theorem, which can be done similarly to Kearns and Singh (2002). There are two sources of failure for the algorithm:
@@
Appendix E More Discussion about experiments
Deterministic Transitions: Deterministic settings are common in many state-of-the-art environments in MORL, particularly when focusing on episodic tasks with short time horizons. These settings are often used to emphasize the core algorithmic contributions without introducing additional complexities from stochastic transitions or long horizons. Nonetheless, we acknowledge that incorporating stochastic environments would better showcase the generality of our approach and highlight this as an important direction for future work.
Model Sizes and Scalability: We recognize that the model sizes used in the experiments are relatively small, primarily due to memory limitations in the tabular setting, and that function approximation would be needed to move to large state space environments as highlighted in our future work.
RAEE Algorithm: RAVI is a model-based approach, and one of the baselines, “Mixture-M” for a model-based mixture policy, uses the model. For the others model-free baselines, we only compare performance at convergence with RAVI to achieve a more fair comparison. To the best of our knowledge, there are no other model-based algorithms in the field specifically designed for optimizing ESR objectives. Due to computational and space constraints, we focused our empirical evaluation on RAVI, as the optimization subroutine is the core algorithmic contribution. That said, we acknowledge this limitation and believe benchmarking RAEE empirically is a worthwhile direction for future research.
Model Accuracy and RAEE: While it is true that RAVI’s performance depends on model accuracy, the central goal of RAEE is precisely to ensure it learns a sufficiently accurate model of the environment to provably guarantee an approximation factor. For future work, comparing RAVI with RAEE would provide valuable insights into the trade-offs between leveraging accurate models and learning them through exploration.